Given the wavefunction of a standing wave: y(x,t) = A sin(πx)cos(2πt), where x and y are in meters and

t is in seconds. The position of the second anti-node from the end x = 0 is at:


1
Expert's answer
2021-12-22T05:42:02-0500

Given:

"y(x,t) = A \sin(πx)\cos(2πt)"


The positions of anti-nodes are given by the equation 


"\sin(\pi x)=\pm 1"



So 


"\pi x=\pi/2+\pi n, \quad n=0,1,2,...\\
x=1/2+n"


The position of second anti-node from origin


"x=1/2+1=3/2\:\rm m=1.5\: m"